OpenStudy (anonymous):

Find the value of tan(a-b) if cos a=-3/5, sin b=5/13, 90

myininaya (myininaya):

Use the difference rule for tan(a-b)

OpenStudy (anonymous):

I'm trying to catch up quickly and didn't get to this lesson. I have no idea what that means.

myininaya (myininaya):

What about the sin difference rule and the cosine difference rule?

OpenStudy (anonymous):

No, I didn't get to any of it. I don't know how to work it.

myininaya (myininaya):

\[\tan(a-b)=\frac{\sin(a-b)}{\cos(a-b)}\] So you don't know the following identities: \[\sin(a \pm b) =\sin(a)\cos(b) \pm \sin(b)\cos(a)\] \[\cos(a \pm b)=\cos(a)\cos(b) \mp \sin(a)\sin(b)\]

myininaya (myininaya):

these are the sum and difference identities for sin and cos

OpenStudy (anonymous):

No, I don't.. Can you please tell me the answer? I'm on a time limit..

myininaya (myininaya):

Oh no. That isn't the way this works.

myininaya (myininaya):

I will help if you are willing to learn and to help yourself.

myininaya (myininaya):

I do not cheat though and OpenStudy does not allow cheating.